# The Unbounded Denominators Conjecture

### v1

@@ 2.2. Equidistribution
(2.2.1) $F(x(\mathbf{t})):=F(x(t_{1}),\ldots,x(t_{d}))\in\mathbf{Z}\llbracket\mathbf{t}\rrbracket\smallsetminus\{0\},$
@@
whose integrality of coefficients follows from Lemma 2.1.2 and our defining assumptions that $p(x)\in\mathbf{Z}\llbracket t\rrbracket$ while all $f_{i}(x(t))\in\mathbf{Z}\llbracket t\rrbracket$. This integrality is the key to the Liouville lower bound. For the Cauchy upper bound, in counterpoint, we would need a pointwise upper bound on the intervening functions $|p(\varphi(\mathbf{z}))^{\mathbf{k}}|$ on the unit polycircle $\mathbf{z}\in\mathbf{T}^{d}$, and here the problem is that while the Monte Carlo heuristic applies on the majority of $\mathbf{T}^{d}$ under $d\to\infty$, with a probability tending to $1$ roughly speaking at a rate exponential in d-d (this follows by Hoeffding’s concentration inequality with (2.2.8) below), the peaks at the biased part of $\mathbf{T}^{d}$ get overwhelmingly large, and a direct extrapolation from (2.2.1) still only leads to a dimension bound with $\sup_{|z|=1}\log{|p\circ\varphi|}$.
@@
To improve the supremum term to the mean term $\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}$, our new idea is to dampen the size at the peaks by firstly multiplying (2.1.3) by a suitably chosen power $V(\mathbf{z})^{M}$ of the Vandermonde polynomial
@@
By applying the Hadamard volume inequality to the Vandermonde determinant in (2.2.2), we recover the following classical result of Fekete, crucial for our approach.
@@
The idea for sifting out the equidistributed tuples $(z_{1},\ldots,z_{d})$ is the following. If the points $z_{1},\ldots,z_{d}$ are poorly distributed in the uniform measure of the circle, the quantity $|V(\mathbf{z})|$ is uniformly exponentially small in $d^{2}$ (Lemma 2.2.9 below). This plays off against the $d^{d/2}=\exp(o(d^{2}))$ bound of Lemma 2.2.3 to sift out the equidistributed points in our pointwise upper bound in the Cauchy integral formula when we extrapolate in § 2.4 below. Liouville’s Diophantine lower bound still succeeds like in André [And04], thanks to the chain rule and the integrality property of coefficients of (2.2.2), but at the Cauchy upper bound we are now aided by the fact that $V(\mathbf{z})^{M}$ is extremely small (an exponential in $Md^{2}$, see (2.2.10)) at the peaks of the pointwise Cauchy bound, where the point $(z_{1},\ldots,z_{d})$ is poorly distributed, while still not too large (subexponential in $Md^{2}$, thanks to Lemma 2.2.3) uniformly throughout the whole polycircle $\mathbf{T}^{d}$.
In the remainder of the current section, we spell out the notion of ‘well-distributed’ and ‘poorly distributed’, and supply the key equidistribution property for the numerical integration step. The following is the standard notion of discrepancy theory.
@@
2.3. Damping the Cauchy estimate
We combine Lemmas 2.2.3 and 2.2.9 for our choice of the damping term $V(\mathbf{z})^{M}$. In the following, all asymptotics are taken under $\alpha\to\infty$ with respect to all other parameters.
@@ 2.3. Damping the Cauchy estimate
2.4. The extrapolation
We can now easily finish the proof of Theorem 2.0.2 by combining the degree estimate (1) of Lemma 2.1.2 with the Cauchy bound (2.3.10) and the integrality properties of the functions $F(x(\mathbf{t}))\in\mathbf{Z}\llbracket\mathbf{t}\rrbracket$ of (2.2.1) and $V(\mathbf{z})\in\mathbf{Z}[\mathbf{z}]$ of (2.2.2).
Let $\beta\geq\alpha$ be the exact order of vanishing of $F(\mathbf{x})$ at the origin $\mathbf{x=0}$. Among the nonvanishing monomials $c\,\mathbf{x^{n}}$ of this minimal order $|\mathbf{n}|=\beta$, choose the one whose degree vector $\mathbf{n}$ has the highest lexicographical ordering. By the chain rule and the minimality of $|\mathbf{n}|$, the normalization condition (2.0.3) on the formal substitution $x(t)$ entails that $c\,\mathbf{t^{n}}$ is a minimal order term in the $t$-expansion $F(x(\mathbf{t}))$. Hence the integrality property (2.2.1) gives that $c\in\mathbf{Z}\smallsetminus\{0\}$ is a nonzero rational integer.

### v2

@@ 2.1. The auxiliary construction
(2.2.1) $H(\mathbf{z}):=h(z_{1})^{D}\cdots h(z_{d})^{D}\cdot F(\varphi(z_{1}),\ldots,\varphi(z_{d}))\in\mathbf{C}\llbracket\mathbf{z}\rrbracket\smallsetminus\{0\}.$
By construction, it vanishes at $\mathbf{z=0}$ to order at least $\alpha$, and it is holomorphic in a neighborhood of the closed unit polydisc because all the split-variables constituents
$\varphi^{*}p;\quad\varphi^{*}f_{1},\ldots,\varphi^{*}f_{m}\in{\mathcal{O}(\overline{D(0,1)})}.$
Let $\beta\geq\alpha$ be the exact order of vanishing of $F(\mathbf{x})\in\mathbf{Q}\llbracket\mathbf{x}\rrbracket\smallsetminus\{0\}$ at $\mathbf{x=0}$, and consider $c\,\mathbf{x^{n}}$ any nonzero monomial of that lowest order $\beta=|\mathbf{n}|$. Since $x(t)\in t+t^{2}\mathbf{Q}\llbracket t\rrbracket$, the term $c\,\mathbf{t^{n}}$ is a lowest order monomial in the formal power series $F(x(\mathbf{t}))\in\mathbf{Z}\llbracket\mathbf{t}\rrbracket$, and so $c\in\mathbf{Z}\smallsetminus\{0\}$. Thus we have the Liouville lower bound:
(2.2.2) $|c|\geq 1.$
On the other hand, (2.2.1) and the normalizations $h(z)\in 1+z\,\mathbf{C}\llbracket z\rrbracket$ and $\varphi(z)\in\varphi^{\prime}(0)z+z^{2}\,\mathbf{C}\llbracket z\rrbracket$ exhibit $c\varphi^{\prime}(0)^{\beta}\,\mathbf{z^{n}}$ as a lowest order monomial in $H(\mathbf{z})$. Since the $\mathbf{z^{n}}$ coefficient is also computed by Cauchy’s integral formula $\int_{\mathbf{T}^{d}}\frac{H(\mathbf{z})}{\mathbf{z^{n}}}\,\mu_{\mathrm{Haar}}(\mathbf{z})$, we have the Cauchy upper bound:
(2.2.3) $|c|\cdot|\varphi^{\prime}(0)|^{\alpha}\leq|c|\cdot|\varphi^{\prime}(0)|^{\beta}\leq\sup_{\mathbf{T}^{d}}|H|.$
To estimate the last supremum under the asymptotic $\alpha\to\infty$ for fixed $d$ and $\kappa$, we note that the sum (2.1.3) is comprised of $(mD)^{d}=\exp\big(o(\alpha)\big)$ terms, all of which are of the form
$\prod_{j=1}^{d}h(z_{j})^{D-k_{j}}\prod_{j=1}^{d}(h(z_{j})\cdot p(\varphi(z_{j})))^{k_{j}}\cdot f_{i_{1}}(\varphi(z_{1}))\cdots f_{i_{d}}(\varphi(z_{d})),$
for some $k_{1},\ldots,k_{d}\in\{0,\ldots,D-1\}$, and with coefficients bounded in magnitude by $\exp\big(\kappa C\alpha+o(\alpha)\big)$.
Every such term is bounded in magnitude on $\mathbf{T}^{d}$ by
$e^{\kappa C\alpha+o(\alpha)}\cdot\max\big\{\sup_{\mathbf{T}}|h|,\,\sup_{\mathbf{T}}|h\cdot\varphi^{*}p|\big\}^{dD}\cdot\max_{1\leq i\leq m}\sup_{\mathbf{T}}|\varphi^{*}f_{i}|^{d}.$
By the triangle inequality, we have in the $\alpha\to\infty$ asymptotic—with respect to a fixed dd—the supremum bound
$\sup_{\mathbf{T}^{d}}\log{|H|}\leq dD\cdot\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}+\kappa C\alpha+o(\alpha)$
$\leq\Big(\kappa C+\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\Big)\cdot\frac{\alpha}{m}+o(\alpha).$
Combining with (2.2.2) and (2.2.3), we get the asymptotic bound
$\alpha\log{|\varphi^{\prime}(0)|}\leq\Big(\kappa C+\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\Big)\cdot\frac{\alpha}{m}\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}+o(\alpha)$
as $\alpha\to\infty$ with respect to the other parameters.
This proves the dimension bound
$m\leq\inf_{\kappa\in(0,1),\,d\in\mathbf{N}}\Big\{\Big(\kappa C+\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\Big)\Big\}\cdot\frac{\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}}{\log{|\varphi^{\prime}(0)|}}.$
Theorem 2.0.8 now follows by firstly letting $d\to\infty$ and then $\kappa\to 0$, and observing that in that limit
$\kappa C+\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\to e,$
by Stirling’s asymptotic and the key point that the constant $C$ depends only on the datum $\{p;\varphi\}$, but not on $d$ nor $\kappa$.
The algebraicity of $f_{i}$ is a fortiori by the finite dimension bound (2.0.3) as all powers of $f_{i}$ satisfy $x^{*}f_{i}^{N}\in\mathbf{Z}[[t]]$ and $\varphi^{*}f_{i}^{N}\in\mathcal{O}(\overline{D(0,1)})$, for any $N\in\mathbf{N}$. ∎
2.3. Canonical factorization and proof of Theorem 2.0.1
At this point Theorem 2.0.1 comes as the immediate combination of Theorem 2.0.8 and the following classical lemma of Nevanlinna.
Lemma 2.3.1 (Nevanlinna [Nev70]).
Every holomorphic function $g:\overline{D(0,1)}\to\mathbf{C}$ admits a quotient representation
$g=\frac{hg}{h},$
where $h:\overline{D(0,1)}\to\mathbf{C}$ is holomorphic with
$h(0)=1\quad\textrm{and}\quad\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|hg|}\big\}\leq\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}.$
Proof.
By a Blaschke product, we are reduced to the case that $g(z)\in\mathcal{O}^{\times}(D(0,1))$ is a functional unit: no zeros or poles inside the unit disc $D(0,1)$. In that case, the function $\log{|g|}$ is harmonic on $D(0,1)$, and the Poisson kernel formula—see (6.1.8) below—together with the canonical decomposition $\log=\log^{+}-\log^{-}$ into positive and negative parts gives the quotient representation
(2.3.2) $g(z)=\frac{\exp\Big(-\int_{\mathbf{T}}\log^{+}{\frac{1}{|g(w)|}}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big)}{\exp\Big(-\int_{\mathbf{T}}\log^{+}{|g(w)|}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big)}.$
In this factorization, the top and bottom both are holomorphic functions of $z\in D(0,1)$, and they both are bounded in absolute value by $\leq 1$, because the Poisson kernel satisfies $\Re\Big(\frac{w+z}{w-z}\Big)>0$ for $1=|w|>|z|$. Furthermore, the bottom in (2.3.2) takes the value $\exp\Big(-\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}\Big)$ at $z=0$.
Thus, in the functional unit case $g\in\mathcal{O}^{\times}(D(0,1))$ and on $D(0,1)$ rather than $\overline{D(0,1)}$, we conclude the proof of the lemma by selecting
$h(z):=\exp\Big(\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}-\int_{\mathbf{T}}\log^{+}{|g(w)|}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big).$
The proof of the general case is completed by the Blaschke product clearance to a functional unit on the open unit disc, and then by shrinking the radius a little bit. The details are, for example, in Nevanlinna [Nev70, §VII.1.4] or Goluzin [Gol69, §VII.5]. ∎
2.4. A first alternative proof
In this section, we complete an idea proposed to us by André as an alternative to our original proof of Theorem 2.0.1 (itself recounted in § 2.5 further down), based on plurisubharmonicity and a lexicographic induction instead of on Cauchy’s formula. We invite the reader at this point to skip ahead directly to §3 on a first pass, as the arithmetic holonomy bound (2.0.7)—the algebraization ingredient that we need for the unbounded denominators conjecture—has already been proved.
2.4.1. Lemma on the lexicographically lowest coefficient
The extrapolation step will now be based on the following analytic lemma, to be applied with $G(\mathbf{z})=F(\varphi(z_{1}),\ldots,\varphi(z_{d}))$, where $F(\mathbf{x})$ is our auxiliary function from Lemma 2.1.2. The lemma reflects the plurisubharmonic property of the multivariable complex functions of the form $\log{|H(\mathbf{z})|}$ with $H(\mathbf{z})$ holomorphic, used inductively on the number of variables $d$.
Lemma 2.4.1.
Consider $G(\mathbf{z})\in\mathbf{C}\llbracket\mathbf{z}\rrbracket\smallsetminus\{0\}$ holomorphic on the closed unit polydisc $\{\mathbf{z}\mid\max_{i=1}^{d}|z_{i}|\leq 1\}$, and let $c\,\mathbf{z^{n}}$ be the lexicographically minimal monomial. Then
(2.4.2) $\log{|c|}\leq\int_{\mathbf{T}^{d}}\log{|G|}\,\mu_{\mathrm{Haar}}.$
Proof.
We induct on the number of variables $d$. For $d=1$, the bound (2.4.2) follows directly from Jensen’s formula, or from the subharmonic property of the function $u(z):=\log{|z^{-n}G(z)|}$, which entails
$\log{|c|}=u(0)\leq\int_{\mathbf{T}}u\,\mu_{\mathrm{Haar}}=\int_{\mathbf{T}}\log{|G|}\,\mu_{\mathrm{Haar}}.$
The last equality uses that the functions $u(z)=\log{|z^{-n}G(z)|}$ and $\log{|G(z)|}$ have the same restriction on the unit circle $\mathbf{T}$.
For the induction step, we write $\mathbf{z}=(z_{1},\mathbf{z}^{\prime})$ and
$G(\mathbf{z})=z_{1}^{n_{1}}H(\mathbf{z}),$
where $H\in\mathbf{C}\llbracket\mathbf{z}\rrbracket$ is holomorphic by our lexicographic minimality assumption. For any fixed $\mathbf{z}^{\prime}\in\mathbf{T}^{d-1}$, by the same argument as the $d=1$ case above, we have
(2.4.3) $\log{|H(0,\mathbf{z}^{\prime})|}\leq\int_{\mathbf{T}}\log{|H(z_{1},\mathbf{z}^{\prime})|}\,\mu_{\mathrm{Haar}}(z_{1})=\int_{\mathbf{T}}\log{|G(z_{1},\mathbf{z}^{\prime})|}\,\mu_{\mathrm{Haar}}(z_{1}).$
By assumption, the lexicographically minimal monomial in $H(0,\mathbf{z}^{\prime})\in\mathbf{C}\llbracket\mathbf{z}^{\prime}\rrbracket$ is $c\,\mathbf{z}^{\prime\mathbf{n}^{\prime}}$, where $\mathbf{n}=(n_{1},\mathbf{n}^{\prime})$. Therefore the induction hypothesis gives
(2.4.4) $\log{|c|}\leq\int_{\mathbf{T}^{d-1}}\log{|H(0,\mathbf{z}^{\prime})|}\,\mu_{\mathrm{Haar}}(\mathbf{z}^{\prime}).$
We complete the induction by integrating the inequality (2.4.3) over $\mathbf{z}^{\prime}\in\mathbf{T}^{d-1}$. ∎
2.4.2. Extrapolation and first alternative proof of Theorem 2.0.1
We apply Lemma 2.4.1 to the $\varphi$-pullback of our $d$-variate auxiliary function:
(2.4.5) $G(z_{1},\ldots,z_{d}):=F(\varphi(z_{1}),\ldots,\varphi(z_{d}))\in\mathbf{C}\llbracket\mathbf{z}\rrbracket\smallsetminus\{0\}.$
This is holomorphic in a neighborhood of the closed unit polydisc, because all the split-variables constituents
$\varphi^{*}p;\quad\varphi^{*}f_{1},\ldots,\varphi^{*}f_{m}\in{\mathcal{O}(\overline{D(0,1)})}.$
Thus, with $c\,\mathbf{z^{n}}$ the lexicographically lowest monomial in $G(\mathbf{z})$, we get from (2.4.2) and Lemma 2.1.2 our Cauchy upper bound:
$\log{|c|}\leq\int_{\mathbf{T}^{d}}\log{|F(\varphi(z_{1}),\ldots,\varphi(z_{d}))|}\,\mu_{\mathrm{Haar}}$
(2.4.6) $\leq dD\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}+\kappa C\,\alpha+o(\alpha),$
asymptotically as $\alpha\to\infty$ with regard to the other parameters. Here, we used the pointwise triangle inequality bound
$\log{|F(x_{1},\ldots,x_{d})|}\leq D\sum_{i=1}^{d}\log^{+}{|p(x_{i})|}+\kappa C\alpha+o(\alpha)$
for $x_{i}:=\varphi(z_{i})$ (note that the sum in (2.1.3) is comprised of $(mD)^{d}=\exp(o(\alpha))$ terms), and integrated this pointwise bound over the unit polycircle $\mathbf{z}\in\mathbf{T}^{d}$.
The Liouville lower bound comes down to the integrality property
(2.4.7) $F(x(t_{1}),\ldots,x(t_{d}))\in\mathbf{Z}\llbracket\mathbf{t}\rrbracket$
inherited from our respective assumptions
$x^{*}p;\quad x^{*}f_{1},\ldots,x^{*}f_{m}\in\mathbf{Z}\llbracket t\rrbracket$
on the split-variables constituents in (2.1.3). By our normalizations $x(t)\in t+t^{2}\mathbf{Q}\llbracket t\rrbracket$ and $\varphi(z)\in\varphi^{\prime}(0)z+z^{2}\mathbf{C}\llbracket z\rrbracket$, the lexicographically lowest term of $G(\mathbf{z})\in\mathbf{C}\llbracket\mathbf{z}\rrbracket$ is equal to $\varphi^{\prime}(0)^{\beta}$ times the lexicographically lowest term of $F(x(\mathbf{t}))\in\mathbf{Z}\llbracket\mathbf{t}\rrbracket$, where $\beta:=|\mathbf{n}|=n_{1}+\cdots+n_{d}\geq\alpha$ is the common total degree of these lexicographically lowest terms in $G(\mathbf{z})$ and $F(x(\mathbf{t}))$. By (2.4.7), this entails that the nonzero coefficient
$c\in\varphi^{\prime}(0)^{\beta}\,\mathbf{Z}\smallsetminus\{0\},$
and hence a fortiori that
(2.4.8) $\log{|c|}\geq\beta\log{|\varphi^{\prime}(0)|}\geq\alpha\log{|\varphi^{\prime}(0)|}.$
We get our requisite dimension bound (2.0.3) on combining the degree bound (1) of Lemma 2.1.2 with the Cauchy upper bound (2.4.6) and the Liouville lower bound (2.4.8), and letting firstly $\alpha\to\infty$, then $d\to\infty$, and finally $\kappa\to 0$.
This completes another proof of Theorem 2.0.1. ∎
2.5. A second alternative proof
The remainder of §2 presents our original argument for Theorem 2.0.1, with the thought that it could still be useful for other settings including potential theory (see 2.5.4). Like § 2.2 and unlike § 2.4, it is based on the leading order jet rather than the overall lexicographically lowest monomial in $F(\mathbf{x})$, and on the pointwise Cauchy integral formula instead of on plurisubharmonicity. Contrastingly to both, it employs a cross-variables equidistribution idea.
2.5.1. Equidistribution
@@
We start out the same way with Lemma 2.1.2, but now aim to extrapolate based directly on the pointwise Cauchy bound. The key idea here is that upon substituting $x_{j}=\varphi(z_{j})$ into (2.1.3), the $d\to\infty$ equidistribution on the circle of the uniform independent and identically distributed points $z_{1},\ldots,z_{d}$ will normally get the constituent monomials in (2.1.3) to grow at most at the integrated exponential rate of $dD\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}$. The problem with directly applying the Cauchy bound as in [And89, VIII 1.6] is that it involves a pointwise upper bound on the intervening functions $|p(\varphi(\mathbf{z}))^{\mathbf{k}}|$ on the unit polycircle $\mathbf{z}\in\mathbf{T}^{d}$, and while the Monte Carlo heuristic applies on the majority of $\mathbf{T}^{d}$ under $d\to\infty$, with a probability tending to $1$ roughly speaking at a rate exponential in d-d (this follows by Hoeffding’s concentration inequality with (2.5.7) below), the peaks at the biased part of $\mathbf{T}^{d}$ get overwhelmingly large, and a direct extrapolation with (2.1.3) in this way still only leads to a dimension bound with $\sup_{|z|=1}\log{|p\circ\varphi|}$.
@@
To improve the supremum term to the mean term $\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}$, we dampen the size at the peaks by firstly multiplying (2.1.3) by a suitably chosen power $V(\mathbf{z})^{M}$ of the Vandermonde polynomial
@@
By applying the Hadamard volume inequality to the Vandermonde determinant in (2.5.1), we recover the following classical result of Fekete, crucial for the present approach.
@@
The idea for sifting out the equidistributed tuples $(z_{1},\ldots,z_{d})$ is the following. If the points $z_{1},\ldots,z_{d}$ are poorly distributed in the uniform measure of the circle, the quantity $|V(\mathbf{z})|$ is uniformly exponentially small in $d^{2}$ (Lemma 2.5.8 below). This plays off against the $d^{d/2}=\exp(o(d^{2}))$ bound of Lemma 2.5.2 to sift out the equidistributed points in our pointwise upper bound in the Cauchy integral formula when we extrapolate in § 2.4.2 below. Liouville’s Diophantine lower bound still succeeds like in André [And04, §5], thanks to the chain rule and the integrality property of coefficients of (2.5.1), but at the Cauchy upper bound we are now aided by the fact that $V(\mathbf{z})^{M}$ is extremely small (an exponential in $Md^{2}$, see (2.5.9)) at the peaks of the pointwise Cauchy bound, where the point $(z_{1},\ldots,z_{d})$ is poorly distributed, while still not too large (subexponential in $Md^{2}$, thanks to Lemma 2.5.2) uniformly throughout the whole polycircle $\mathbf{T}^{d}$.
In the remainder of the current subsection, we spell out the notion of ‘well-distributed’ and ‘poorly distributed’, and supply the key equidistribution property for the numerical integration step. The following is the standard notion of discrepancy theory.
@@
2.5.2. Damping the Cauchy estimate
We combine Lemmas 2.5.2 and 2.5.8 for our choice of the damping term $V(\mathbf{z})^{M}$. In the following, all asymptotics are taken under $\alpha\to\infty$ with respect to all other parameters.
@@ 2.5.2. Damping the Cauchy estimate
2.5.3. The extrapolation
Finally we combine the degree estimate (1) of Lemma 2.1.2 with the Cauchy bound (2.5.22) and the integrality properties of the functions $F(x(\mathbf{t}))\in\mathbf{Z}\llbracket\mathbf{t}\rrbracket$ of (2.1.3) and $V(\mathbf{z})\in\mathbf{Z}[\mathbf{z}]$ of (2.5.1).
Let $\beta\geq\alpha$ be the exact order of vanishing of $F(\mathbf{x})$ at the origin $\mathbf{x=0}$. Among the nonvanishing monomials $c\,\mathbf{x^{n}}$ of this minimal order $|\mathbf{n}|=\beta$, choose the one whose degree vector $\mathbf{n}$ has the highest lexicographical ordering. By the chain rule and the minimality of $|\mathbf{n}|$, the normalization condition (2.0.2) on the formal substitution $x(t)$ entails that $c\,\mathbf{t^{n}}$ is a minimal order term in the $t$-expansion $F(x(\mathbf{t}))$. Hence the integrality $f(x(\mathbf{t}))\in\mathbf{Z}\llbracket\mathbf{t}\rrbracket$ gives that $c\in\mathbf{Z}\smallsetminus\{0\}$ is a nonzero rational integer.